1. Find b. Round to the nearest whole inch.
the triangle ABC with length BC 14 in. Angle B = 82 degrees and Angle A= 54 degrees. And length AC = b
(Points : 1)
17 in.
6 in.
11 in.
14 in.

Answer by Alan3354(69443)

the triangle ABC with length BC 14 in. Angle B = 82 degrees and Angle A= 54 degrees. And length AC = b
-----------------
BC = a
-----------
a/sin(A) = b/sin(B)
14/sin(54) = b/sin(82)
b = 14*sin(82)/sin(54)
b =~ 17 in